OpenStudy (anonymous):

Iogistics it says log7^343 the seven is small and it says to Evaluate each expression

OpenStudy (anonymous):

\[ \log_7{343} \]??

OpenStudy (anonymous):

since \(343=7^3\) you should get the answer instantly

OpenStudy (anonymous):

The first thing you should do is factor the expression out.

OpenStudy (anonymous):

I don't know how to do this i just moved schools and they are ahead he told me to solve for X the example says

OpenStudy (anonymous):

When ever you have a logarithm, you want to factor the number... do you know how to do that?

OpenStudy (anonymous):

the answer to \(\log_2(32)\) is 5, because \(2^5=32\)

OpenStudy (anonymous):

\[\log_b(x)=y\iff b^y=x\] so if you have \[\log_2(32)=y\] it is the same as solving \[2^y=32\] for \(y\)

OpenStudy (anonymous):

no, today was the first day I've been introduced to this and i understood the rest of the worksheet i just don't know how to figure out x on this part..

OpenStudy (anonymous):

since \(2^5=32\) we know \(\log_2(32)=5\) and since \(7^2=343\) we have \(\log_7(343)=3\)

OpenStudy (anonymous):

sorry typo on the last line, i meant to say since \(7^3=343\) then \(\log_7(343)=3\)

OpenStudy (anonymous):

oh, so find the little number that will multiply the other number to equal the larger number

OpenStudy (anonymous):

it is the exponent

OpenStudy (anonymous):

>.< sorry exponent thank you i believe i get it now